### v1

# DRAGON: LLM-Driven Decomposition and Reconstruction Agents for Large-Scale Combinatorial Optimization

Abstract.
Large Language Models (LLMs) have recently shown promise in addressing combinatorial optimization problems (COPs) through prompt-based strategies. However, their scalability and generalization remain limited, and their effectiveness diminishes as problem size increases, particularly in routing problems involving more than 30 nodes. We propose DRAGON, which stands for Decomposition and Reconstruction Agents Guided OptimizatioN, a novel framework that combines the strengths of metaheuristic design and LLM reasoning. Starting from an initial global solution, DRAGON autonomously identifies regions with high optimization potential and strategically decompose large-scale COPs into manageable subproblems. Each subproblem is then reformulated as a concise, localized optimization task and solved through targeted LLM prompting guided by accumulated experiences. Finally, the locally optimized solutions are systematically reintegrated into the original global context to yield a significantly improved overall outcome. By continuously interacting with the optimization environment and leveraging an adaptive experience memory, the agents iteratively learn from feedback, effectively coupling symbolic reasoning with heuristic search. Empirical results show that, unlike existing LLM-based solvers limited to small-scale instances, DRAGON consistently produces feasible solutions on TSPLIB, CVRPLIB, and Weibull-5k bin packing benchmarks, and achieves near-optimal results (0.16% gap) on knapsack problems with over 3M variables. This work shows the potential of feedback-driven language agents as a new paradigm for generalizable and interpretable large-scale optimization.
Key words and phrases:
Combinatorial Optimization, Large-scale, Metaheuristics
Zhuoyi Lin is the corresponding author.
1. Introduction
Figure 1. Illustration of DRAGON on a TSP instance, where metadata represent the coordinates of locations to visit, serving as the main environment for the agents. The decomposer identifies a suboptimal active segment {4,6,7,8} (grayed) from a global solution. The segment is locally refined and reintegrated by the reconstructor to improve the global tour.
Combinatorial optimization problems (COPs), such as the Traveling Salesman Problem (TSP), Vehicle Routing Problem (VRP), and Knapsack Problem (KP), are notoriously challenging due to their NP-hard nature Ausiello et al. (2012). These COPs are typically addressed using exact algorithms or heuristic methods, which often require substantial manual design and parameter tuning Rokbani et al. (2021); Cacchiani et al. (2022); Ma et al. (2023); Lin et al. (2024); Feng et al. (2025), thereby limiting their scalability and adaptability to varying problem instances and domains.
Large Language Models (LLMs) have led to a series of breakthroughs for various challenging problems in natural language processing and artificial intelligence, including language understanding, text generation, code synthesis, and reasoning tasks (Brown et al., 2020; Chowdhery et al., 2023; Achiam et al., 2024). Recent evidence reveals that LLMs have emerged as powerful tools capable of addressing a variety of COPs, leveraging their reasoning capabilities for direct solution generation Yang et al. (2024); Wang et al. (2024) or heuristic design Ye et al. (2025); Romera-Paredes et al. (2024); Liu et al. (2024b); Zhu et al. (2025). Their inherent strengths in abstraction, generalization, and semantic understanding allow them to generate promising solutions even without explicit algorithmic instructions. However, despite these advances, the current capabilities of LLMs for directly generating solutions remain largely confined to relatively small-scale instances, such as the TSP with fewer than 30 nodes Yang et al. (2024); Iklassov et al. (2024); Liu et al. (2024a). As problem size and complexity grow, LLM-based solutions often deteriorate due to inherent limitations, including restricted context length, reduced logical coherence, and difficulty in representing combinatorial structures Iklassov et al. (2024); Zhao et al. (2025). These limitations significantly impede the practical deployment of LLM-driven methods in realistic large-scale applications, such as logistics, transportation, and supply chain management, where problems commonly involve hundreds to thousands of nodes Bengio et al. (2021).
In this study, we aim to answer two research questions: (1) Can an LLM agent, without solving the full COP, identify and isolate regions of the solution that are likely suboptimal or have improvement potential? (2) Can an LLM agent effectively solve small-scale COPs locally while following additional, customized non-trivial constraints? These questions highlight the core challenges of our framework: utilizing LLMs to guide decomposition by detecting potential improvement areas, and ensuring reconstructed solutions remain feasible.
Prompt-based LLM methods often fail on large-scale instances due to the absence of domain knowledge and the context length limitation, while code-generation approaches require extensive task-specific training to develop robust heuristics, leading to substantial computational costs and poor cross-domain transferability. Meanwhile, classical metaheuristics such as divide-and-conquer and large neighborhood search (LNS) Pisinger and Ropke (2018) exhibit strong scalability in large-scale COPs by iteratively decomposing and refining solutions, though they rely heavily on handcrafted heuristics and expert knowledge. To address the gap, we propose DRAGON, a framework that transforms complex large COP solving into decomposition and reconstruction tasks guided by LLM agents. DRAGON first strategically identifies regions with high potential of improvement, decomposing large-scale COPs into context-manageable subproblems, and locally optimizes each subregion through reconstruction.
To concretely illustrate this idea, Figure 1 demonstrates an example of how the proposed Dragon improves TSP solution locally and globally. Given an initial global solution, the decomposition agent $\mathcal{D}$ identifies nodes $\{4,6,7,8\}$ (highlighted in gray) as a subproblem with high potential for further improvement, while the remaining nodes remain static and impose boundary constraints. During reconstruction, the suboptimal local segment (4-8-6-7) is locally optimized with LLM agents, leading to a better sub-solution (4-6-7-8). Consequently, the improved local segment is integrated into the global solution, yielding a reduced tour length solution.
Overall, our contributions are threefold: (1) To the best of our knowledge, this is the first work to demonstrate that LLM agents can be effectively leveraged to directly generate high-quality solutions for large-scale COPs, opening new possibilities for LLM-driven optimization. (2) We propose DRAGON, a divide-and-conquer framework that dynamically decomposes large-scale COPs and subsequently refines compact subproblems by state passing between agents. (3) We empirically validate DRAGON on large-scale COPs benchmarks, demonstrating substantial improvements in solution quality and scalability compared to state-of-the-art LLM-based baselines. Our findings validate the feasibility of leveraging LLM agents for large-scale COPs that are prevalent in real-world applications.
2. Related Work
LLMs for Combinatorial Optimization. Recent studies have increasingly examined the capability of LLMs to solve COPs by leveraging prompt engineering techniques. Among these efforts, a prominent direction involves using LLMs directly as solution generators, where carefully designed prompts and enriched input information guide the models to produce viable solutions. Early research by Wang et al. (2024) provided evidence that LLMs could solve certain graph-based COPs, although this capability was strongly tied to the quality of the provided prompts. Subsequent work has sought to refine this process by incorporating contextual information. Examples include integrating graph structural and topological data Wasserkrug et al. (2024), existing heuristic solutions Iklassov et al. (2024), explicit relationships between solutions and objectives Yang et al. (2024), and visual representations of problems or solutions Huang et al. (2024); Elhenawy et al. (2024a). Another research pathway involves leveraging LLMs as components within structured algorithms or frameworks. For instance, Liu et al. (2024a) demonstrated the utility of LLMs in evolutionary computation by prompting them to perform essential algorithmic operations such as selection, crossover, and mutation. Similarly, Elhenawy et al. (2024b) deployed multiple LLMs as collaborative agents, each responsible for distinct roles such as initialization, critique, and evaluation within a predefined optimization pipeline. Despite these promising advancements, current capabilities for directly generating solutions via LLMs remain largely restricted to relatively small-scale problems, typically limited to instances such as TSP with fewer than 30 nodes. To circumvent this limitation, Iklassov et al. (2024) introduced a decomposition approach wherein an LLM assesses problem complexity autonomously. If deemed difficult, the model recursively subdivides the problem into simpler, manageable subproblems. Nonetheless, its scalability remains limited.
Beyond direct solution generation, other studies have explored leveraging LLMs for heuristic design Romera-Paredes et al. (2024); Liu et al. (2024b); Ye et al. (2025) and mathematical modeling Xiao et al. (2023); AhmadiTeshnizi et al. (2024); Jiang et al. (2025b). The former aims to use LLMs to discover heuristics expressed in code that have potential to outperform those crafted by human experts, while the latter focuses on translating natural language problem descriptions into mathematical formulations compatible with traditional OR solvers. More recently, research has investigated training large language models for end-to-end combinatorial optimization, rather than relying solely on prompting paradigms Jiang et al. (2025a, 2024). These directions fall outside the scope of this work, and we refer interested readers to Da Ros et al. (2025) for a comprehensive survey.
Large-Scale Combinatorial Optimization. Another research direction related to this work aiming to design metaheuristics for solving large-scale COPs, either with machine learning Li et al. (2021); Sun and Yang (2023) or through traditional methods Shaw (1998); Arnold et al. (2019). These approaches generally follow decomposition or divide-and-conquer principles. While they have shown promising results, their effectiveness often relies on problem-specific strategies (e.g., domain-specific decomposition rules), which limits their generality. In this work, we propose DRAGON, a general decomposition-reconstruction framework leveraging the semantic capabilities of LLMs. Unlike previous methods, DRAGON identifies suboptimal regions without explicit domain-specific rules and iteratively reconstructs globally consistent solutions. Our framework effectively addresses LLM context-length constraints, combining learned decomposition advantages with the generalization provided by LLM agents.
3. Methodology
Figure 2. Overview of the state passing among agents in DRAGON framework. With (a) given COP data input and current global solution, the DRAGON pipeline process alternates between two key stages: (b) A decomposition step uses an LLM to split a large-scale COP into manageable active and static segments. These are compressed into a smaller subproblem; (c) A reconstruction step solves the reduced problem with additional constraints to yield a refined local solution. This is then reintegrated into the global solution.
DRAGON is a two-stage framework that enables LLM agents to effectively solve large-scale COPs via state passing communication. As illustrated in Figure 2, DRAGON iteratively alternates between two core tasks: Decomposition and Reconstruction. Starting from an initial solution generated by any fast heuristic, DRAGON progressively refines it by identifying subregions with high improvement potential and solving localized subproblems under explicit constraint awareness. In each iteration, the current solution is decomposed into manageable subproblems, locally optimized through reconstruction agents, and then reassembled into a globally improved solution. This hierarchical process unifies local and global optimization, allowing DRAGON to scale from small instances to complex large-scale COPs. The example workflow illustrated in Figure 1 is summarized in Algorithm 1, where implicit communication between agents occurs via solution state passing, and the subsequent sections detail each component of the framework.
Algorithm 1 DRAGON for Large-Scale COPs
0:  Metadata $M$, initial solution $S_{0}$, maximum time $T_{\max}$, rejection threshold NN
0:  Improved solution $S$
1:  $S\leftarrow S_{0}$ // Current solution
2:  $t\leftarrow 0$ // Time counter
3:  $r\leftarrow 0$ // Rejection counter
4:  while $t<T_{\max}$ and $r<N$ do
5:   $(\mathbf{a},\mathbf{s})\leftarrow\mathcal{D}(M,S)$ // Decomposition
6:   $(M^{\prime},S^{\prime},C)\leftarrow\texttt{Compress}(M,\mathbf{a},\mathbf{s})$
7:   $S^{\prime}_{\text{new}}\leftarrow\mathcal{R}(M^{\prime},S^{\prime},C)$ // Reconstruction
[…]
9:    $S^{\prime}\leftarrow S^{\prime}_{\text{new}}$
10:    $S\leftarrow\texttt{Integration}(S^{\prime},S)$
11:    $r\leftarrow 0$ // Reset rejection counter
12:   else
13:    $r\leftarrow r+1$
14:   end if
15:   $t\leftarrow\texttt{ElapsedTime}()$ // Reach running time limit.
16:  end while
17:  return $S$
3.1. Framework Inputs and Solution Updates
DRAGON interacts with problem environment that contains two fundamental inputs: metadata $M$ and the current solution $S$. Metadata $M$, represented as structured JSON, encodes essential problem-specific parameters, and the solution $S$ is represented as an ordered sequence of elements, such as city indices for TSP or item identifiers for Knapsack, and is formatted in either JSON or XML, as illustrated in Figure 2(a).
Formally, given metadata $M$, the global solution at iteration $i$ is obtained as follows:
$S_{i}=\begin{cases}\texttt{Fast\_Heuristic}(M),&i=0\\ \texttt{Integration}(S^{\prime}_{i-1},S_{i-1}),&i>0\\ S^{\prime}_{i-1}=\mathcal{R}\left(\mathcal{D}\left(M,S_{i-1}\right)\right),&i>0\end{cases}$ (1)
Here, $S^{\prime}_{i-1}$ denotes the locally refined sub-solution from the previous iteration, identified by the decomposer $\mathcal{D}$ and enhanced by the reconstructor $\mathcal{R}$ (both detailed in subsequent sections). Integration involves simply concatenation and sorting operations, while specific implementation of Fast_Heuristic are provided in the experimental section. Furthermore, a new solution is conditionally accepted using a probabilistic criterion (assuming minimization):
$\texttt{Accept}(S_{1},S_{2})=\min\left\{1,\exp\left(-\frac{f(S_{2})-f(S_{1})}{T}\right)\right\}$ (2)
where $f$ denotes the objective function and $T$ is a temperature parameter controlling the acceptance probability. As shown in Algorithm 1, we substitute the current local solution $S^{\prime}_{i}$ and its $\mathcal{R}$ processed version $S^{\prime}_{\text{new},i}$ into Eq. 2 as $S_{1}$ and $S_{2}$, respectively, and accepted solutions are then integrated back into the global solution to ensure consistency and feasibility. It is worth noting that even when the new local solution is worse i.e., $f(S^{\prime}_{\text{new},i})\geq f(S^{\prime}_{i})$, there remains a small chance of acceptance and integration, allowing DRAGON to occasionally accept inferior solutions, helping it escape local optima, promote exploration of the solution space, and maintain a balancing between exploitation and exploration.
Balancing exploitation (greedy improvement) and exploration (diversification) is critical to the performance of heuristic search methods. In many cases, domain experts must carefully design problem-specific strategies to achieve this balance, and the convergence rate of the solution gap can vary significantly depending on the designer’s expertise. However, for large-scale COPs where analyzing the problem structure is extremely challenging and generalization across domains is difficult, we leverage LLM-based agents, which are well-suited for both exploitation and exploration, because they can exploit stored experiences from previous trails, while their generative and adaptive capabilities enable diverse solution proposals, supporting broader exploration of the solution space.
3.2. Decomposition for Large-Scale COPs
A central challenge in applying LLMs to large-scale COP decomposition lies in their limited capability to reason over complex global constraints. Although existing studies demonstrate the potential of LLMs in solving COPs Wilson et al. (2023); Yang et al. (2024), their capability to accurately detect and isolate sub-problems suitable for local refinement within large-scale COPs remains largely unexplored. To this end, we propose a decomposition agent ($\mathcal{D}$) which leverages LLMs to strategically partition COPs into smaller, manageable sub-problems for further refinement, rather than directly solving the entire COP. Specifically, the decomposer $\mathcal{D}$ functions as a high-level planner, analyzing metadata $M$ and the current solution $S_{i}$. It partitions the solution into active segments ($\mathbf{a}_{i}$), which have substantial improvement potential, and static segments ($\mathbf{s}_{i}$), which remain unchanged during the current iteration. Specifically, decomposition can be formulated as:
$(\mathbf{a}_{i},\mathbf{s}_{i})=\mathcal{D}(M,S_{i})$ (3)
DRAGON demonstrates that well-designed prompts (illustrated in Figure 2(b) and Supplementary Materials (Prompt designs)) enable LLMs to effectively identify segments with significant potential for improvement while introducing additional constraints ($C_{i}$) to maintain feasibility and global consistency:
$(M^{\prime}_{i},S^{\prime}_{i},C_{i})=\texttt{Compress}(M,\mathbf{a}_{i},\mathbf{s}_{i})$ (4)
where $M^{\prime}_{i}\subseteq M$ denotes the subset of original metadata relevant only to the active segments. For example, in the TSP scenario illustrated in Figure 1, the decomposer identifies an active segment $\mathbf{a}_{i}=\{4,6,7,8\}$ and a static segment $\mathbf{s}_{i}=[9,0,1,2,3,5]$. The Compress step significantly reduces the size of the local data $|M_{i}^{\prime}|<<|M|$, making it manageable to downstream agents. However, it will introduce extra specific constraints, which are a set of compressed representations of static segment $\mathbf{s}_{i}$ (e.g., enforce edge preservation like $(9,5)\in C_{i}$).
3.3. Reconstruction with Constraint Integration
Recent studies Yang et al. (2024); Liu et al. (2024a); Iklassov et al. (2024) highlight the effectiveness of LLMs in solving small-scale COPs, typically with fewer than 30 decision variables. Building on this capability, our reconstruction agent ($\mathcal{R}$) is designed to solve compressed COP instances iteratively, derived from the decomposition stage.
Figure 3. Constraint satisfaction in reconstruction is ensured via experience accumulation.
Specifically, reconstructor $\mathcal{R}$ utilizes localized metadata $M^{\prime}_{i}$ augmented with explicit constraints $C_{i}$ to ensure global feasibility and consistency. Constraints $C_{i}$ are explicitly integrated into natural language within the LLM prompt, with the formats varying by COP type. Formally, the reconstruction process is defined as:
$S^{\prime}_{\text{new},i}=\mathcal{R}(M^{\prime}_{i},S^{\prime}_{i},C_{i})$ (5)
where $M^{\prime}_{i}$ encapsulates localized metadata, and $C_{i}$ encodes global consistency requirements derived from static segments $\mathbf{s}_{i}$.
Note that the format and interpretation of constraints depend on the specific COP type. In routing problems such as TSP and CVRP, constraints typically enforce particular edge sequences to ensure route continuity, while in packing problems like MKP or BPP, constraints explicitly mandate item inclusion or exclusion. For instance, a constraint prompt for TSP may state: “the following edges must be visited consecutively, with no additional points permitted between them.” It is essential to maintain consistency between constraints and compressed metadata. For example, required edges $(a,b)$ in CVRP represent condensed segments from the original route, with associated demands accurately reflecting the cumulative demands of intermediate nodes.
As illustrated in Figure 3, reconstruction agent utilizes iterative experience storage collected along previous trails to self-revise candidate solutions until all specified constraints are satisfied. The agent’s experience storage records all previous attempts, regardless of whether they result in feasible solutions or not, along with annotations on the type and cause of any constraint violations. This memory allows the agent to learn from past infeasible solutions, avoid repeating similar errors, and identify patterns within feasible solutions to guide future improvements. Through this process of iterative refinement, localized solutions are gradually adjusted to remain coherent and globally feasible. Extensive experiments detailed in the experimental section and Supplementary Materials demonstrate that, when guided by well-structured prompts and clearly defined constraints, LLM agents can consistently generate high-quality solutions to constrained subproblems.
4. Experiments
4.1. Setup
4.1.1. Problem Descriptions
To evaluate the effectiveness and generalizability of the proposed DRAGON framework, we conduct experiments on four representative COPs across routing, packing and assignment domains:
  • Traveling Salesman Problem (TSP): A classical combinatorial routing problem that seeks the shortest possible tour visiting each city exactly once and returning to the starting point.
  • Capacitated Vehicle Routing Problem (CVRP): Extension of TSP with multiple vehicles and demand constraints. The goal is to minimize total travel distance while satisfying delivery demands.
  • Bin Packing Problem (BPP): Classical combinatorial optimization problem where items of various sizes must be packed into a minimal number of fixed-capacity bins without exceeding their limits.
  • Multiple Knapsack Problem (MKP): A generalization of the 0–1 knapsack problem involving multiple independent knapsacks. Unlike the multi-dimensional knapsack problem Cacchiani et al. (2022), where each item is constrained by several resource limits, the MKP aims to maximize the total value by optimally assigning items to knapsacks without exceeding their individual capacities.
4.1.2. Datasets
  • TSP We use 77 benchmark instances from TSPLIB Reinelt (1991) of the EUC_2D type, with sizes ranging from 50 to 20k nodes, covering a broad range of scales and spatial layouts. Oracle: optimal values or lower bounds reported in Reinelt (1991).
  • CVRP We select subsets of 19 instances from CVRPLIB, including both X-type Uchoa et al. (2017) (up to 1000 nodes) and larger XML-type instances Queiroga et al. (2022) distribution, which was synthesized at large scale (up to 5k nodes) by Zhou et al. (2023). Oracle: optimal values or lower bounds reported in Uchoa et al. (2017); Zhou et al. (2023).
  • BPP We adopt instances from FunSearch Romera-Paredes et al. (2024) and EoH Liu et al. (2024b) of all Weibull-5k, which is reflecting real-world scheduling and allocation scenarios. Oracle: L2 lower bounds determined by Romera-Paredes et al. (2024).
  • MKP Following Google OR-Tools guidelines Perron and Furnon (2025), we generate 10 synthetic instances by with knapsack capacities uniformly sampled from $U(100,500)$, using 10 to 100 knapsacks. Item values and weights drawn uniformly from $U(1,100)$. All instances will be released in supplementary JSON files. Oracle: optimal values or uppers bound searched by solver.
4.1.3. Implementation Details
We evaluate our framework using four representative LLMs accessed via API: two general-purpose models, OpenAI’s gpt-4o and gpt-4.1, and two reasoning-specialized models, OpenAI’s o3 and DeepSeek’s r1. Due to time and cost constraints, we limit our experiments to these models. Prompting strategies for each stage of DRAGON are illustrated in Figure 2, with complete prompt templates provided in the Appendix (Prompt designs). We adopt simple global solution initialization algorithms to each problem: 1) Random Insertion Azar (1994) for TSP, 2) Greedy Nearest Neighbor heuristic Mohammed et al. (2017) for CVRP, and 3) First-Fit Decreasing Baker (1985) for both BPP and MKP. For comparative evaluation, we implement OR-Tools Perron and Furnon (2025) solvers as: CVRP is solved via the routing model Furnon and Perron (2025) with Guided_Local_Search [^1] https://developers.google.com/optimization/routing/routing_options as the metaheuristic, while BPP uses the CP-SAT Perron and Didier (2025). All experiments run on an Ubuntu 20.04 server with an Intel Core i9-10900X (20 cores at 4.6 GHz) without GPU. For detailed prompt designs used in the decomposition and reconstruction processes, please refer to the Supplementary Materials.
4.1.4. Feasibility check
Feasibility checking for solutions obtained by any method is straightforward. We implement a checker based on the inherent constraints of each COP type. If a solution is found to be infeasible, it is stored in the agent’s memory along with comments explaining the reason for infeasibility. This feedback helps the agent avoid repeating the same mistakes, thereby reducing search time and accelerating convergence to feasible solutions. In the results presented later, we denote an infeasible solution as “infe” and use “inf” to indicate cases where the solving time exceeds the given time limit $T_{\max}$.
4.1.5. Metrics
We evaluate performance using four metrics: (1) Optimality Gap defined as $Gap=\frac{|v-v^{*}|}{v^{*}}\times 100\%$, where $v$ and $v^{*}$ are the objective value and the optimal value, respectively; (2) Running Time (seconds); (3) Input/Output Token Counts; and (4) Number of API Calls. All experiments use consistent settings with a 1-hour time limit $T_{\max}=3600$ and early stopping after $N=5$ consecutive non-improving iterations.
4.2. Performance Evaluation
Table 1. Average optimality gap (%) on routing problems using subsets from TSPLIB (EUC_2D) and CVRPLIB (Set X/XML).
Method TSPLIB (EUC_2D)
(size, k 0.05-0.5 0.5-1 1-2 2-5 5-20 All
#) 42 6 15 7 7 77
OPRO inf inf inf inf inf inf
SGE infe infe infe infe infe infe
LMEA 820.30 2279.35 3578.21 inf inf inf
ReEvo(c) 12.69 16.43 17.01 inf inf inf
ReEvo(a) 8.17 16.61 18.23 21.20 18.62 13.10
DRAGON 9.74 11.24 15.24 19.37 15.42 12.24
Method CVRPLIB (X/XML)
(size, k 0.1-0.2 0.2-0.5 0.5-1 1-2 2-5 $\geq$5 All
#) 2 3 5 3 3 3 19
OR-Tools inf inf inf inf inf inf inf
ReEvo(a) 21.49 19.27 29.44 25.39 22.00 11.24 24.17
DRAGON 25.56 29.35 26.73 15.51 15.48 6.45 20.15
To evaluate performance on large-scale routing problems (TSP and CVRP), Table 1 presents the average optimality gap (%) across different instance size groups, comparing DRAGON with existing LLM-based solvers, including prompt-based methods (OPRO Yang et al. (2024), SGE Iklassov et al. (2024)) and code-generation approaches (LMEA Liu et al. (2024a), ReEvo Ye et al. (2025), with ReEvo(a) using Ant Colony Optimization (ACO) and ReEvo(c) employing a constructive heuristic). To ensure a fair comparison, all LLM-based methods use gpt-4o, with token usage and inference time recorded. Instances are grouped by node size for conciseness, and detailed results are listed in Supplementary Materials.
OPRO fails to return valid solutions for all groups, as some hard instances in groups exceed the runtime limit, resulting in an infinite gap inf. SGE consistently produces infeasible solutions (infe), likely because it lacks explicit feasibility checks, since its original design is for small problems (under 30 nodes). LMEA handles slightly larger instances but exhibits high gaps and fails beyond 2k-node due to prompt length limits and generation timeouts.
Figure 4. Average Optimality Gap (%) (log scale) across problem size groups for different methods on TSPLIB.
As shown in Figure 4, results on TSPLIB reveal that SGE is excluded due to infeasible solutions, while OPRO is omitted as “inf” that no solution found due to its slow inference time across all groups. LMEA is able to solve up to medium-sized instances but exhibits relatively large optimality gaps. ReEvo(c) achieves strong results up to the medium-size group, whereas ReEvo(a) excels on smaller instances. Although DRAGON does not outperform across every group, it demonstrates clear superiority on larger-scale problems, benefiting from its task allocation mechanism, where decomposition and reconstruction collaborate effectively.
Figure 5. Optimality Gap (%) for different methods on CVRPLIB.
Figure 5 presents the methods capable of obtaining feasible solutions for CVRP instances. As CVRP is inherently more complex than TSP, DRAGON outperforms ReEvo(a) primarily on large-scale problems, where its decomposition–reconstruction mechanism demonstrates greater advantages.
These findings highlight a major limitation of pure prompt-based methods: while capable of solving small to medium-scale COPs, their performance degrade significantly as problem size increases. This is mainly due to long context lengths and high decoding cost—making token-by-token generation is unstable when encounter large input loads. Code-generation methods like ReEvo improve scalability. We tested ReEvo’s released heuristic code, including both constructive (ReEvo(c)) and ACO-enhanced (ReEvo(a)) variants. On TSPLIB, DRAGON consistently outperforms ReEvo(c) across all size groups and also surpasses ReEvo(a) on all groups beyond 500, and achieving an average gap reduction of 0.86%. On CVRPLIB, where OR-Tools fails to find feasible solutions within time limit, DRAGON leads around 3.98% lower gaps than ReEvo(a) on all cases. Overall, DRAGON demonstrates the best average performance across routing benchmarks, validating its effectiveness and scalability.
4.3. Ablation Study
4.3.1. Strategies for Decomposition and Reconstruction
Table 2. Domain expertise levels across decomposition–reconstruction strategy combinations. Each cell shows expertise needed: ✗ (none), ✓ (some), ✓✓ (extensive) for decomposition and reconstruction, respectively.
Decomposition Reconstruction
Heuristic Solver LLM
Random (✗, ✓) (✗, ✓✓) (✗, ✗)
Heuristic (✓, ✓) (✓, ✓✓) (✓, ✗)
LLM (✗, ✓) (✗, ✓✓) (✗, ✗)
Figure 6. Ablation study of DRAGON across 9 decomposition–reconstruction combinations as listed in Table 2. Bars show the optimality gap (%) compared to the known optimal value on a TSPLIB subset.
DRAGON supports configurable strategies for each task, allowing us to evaluate the impact of different implementations. Table 2 presents a $3\times 3$ combination of decomposition and reconstruction strategies, along with the domain expertise required for each pair, offering insight into their implementation difficulty. For decomposition, “Random” selects elements arbitrarily, “Heuristic” applies domain-specific rules, and “LLM” leverages large language models. Similarly for strategies in reconstruction, the only new word “Solver” refers to implement general purpose solver. Figure 6 shows the performance of DRAGON on a TSPLIB subset using these nine combinations.
[…]
4.3.2. Impact of LLM backbones
Figure 7. Comparison of different LLM models by: (a) optimality gap, (b) number of API calls, (c) input tokens, (d) output tokens, (e) running time. We evaluate general-purpose models (gpt-4o, gpt-4.1) and reasoning-oriented models (o3, r1). Gray bars indicate incomplete results due to token limit violations on the largest instance.
As shown in Figure 7, we evaluate the impact of different LLM backbones within the DRAGON framework, including general-purpose models (OpenAI gpt-4o, gpt-4.1) and reasoning models (OpenAI o3, DeepSeek r1). Due to cost considerations, we do not evaluate all TSPLIB instances. As r1 exceeds the model’s input token limit of 65,536 on instances larger than 10k nodes, it fails to produce responses in the largest case (listed in Appendix (Experiment results)).
While o3 achieves the lowest average optimality gap, it incurs extremely high output token counts. Given that OpenAI’s API typically prices output tokens at 4 times the rate of input tokens, this leads to significantly higher costs and longer inference time. r1 exhibits a relatively low number of API calls, suggesting that each call takes longer—likely due to more intensive reasoning per iteration. Considering the trade-off between objective quality, inference speed, and token efficiency, gpt-4.1 achieves the best overall balance and emerges as the current most practical choice.
4.4. Generalization Case Study
To assess the robustness of DRAGON, we evaluate its performance across more domains, it highlights the method’s ability to generalization for the subsequent results.
4.4.1. Bin Packing Problem
Table 3. Average optimality gap (%) w.r.t. L2 lower bound and inference time (sec.) on the Weibull-5K dataset.
Method ReEvo(a) FunSearch EoH EoH expert Ours
Gap (%) $\downarrow$ 3.46 0.69 0.66 0.55 0.33
Time (sec.) $\downarrow$ 56.677 2.292 - - 487.873
In addition to routing problems, we evaluate DRAGON on the packing domain. As shown in Table 3, DRAGON consistently outperforms ReEvo(a), FunSearch, EoH, and EoH expert on the Weibull-5K dataset. We omit OR-Tools solvers (both MP and CP-SAT backends) as they failed to produce feasible solutions before timeouts.
Compared to routing, packing problems are generally easier for LLMs to reason about due to their more intuitive and straightforward constraints. For example, merging static segments ($\mathbf{s}$) into packed items requires less complex reasoning, allowing DRAGON to scale more effectively to larger instances. Results for EoH and EoH expert are reported directly from Liu et al. (2024b), where running time was not provided. While DRAGON achieves the lowest optimality gap, it is slower than code-generation based methods. However, code-generation approaches come with high API costs and substantial time and resources to evolve or fine-tune high-quality heuristic functions. These overheads are not captured in current comparison. Each paradigm has its strengths. For large batches of similar instances, code-generation methods are effective as the learned heuristic can be reused across the dataset. In contrast, prompt-based approaches like DRAGON are better suited for dynamic environments where instance structures vary, offering flexibility without the need for retraining or code evolution.
4.4.2. Multiple Knapsack Problem
Figure 8. Comparison on synthetic Multiple Knapsack Problem instances between DRAGON and OR-Tools (CP-SAT).
[…]
As shown in Figure 8, CP-SAT typically achieves lower optimality gaps and can even find optimal solutions when the instance size is manageable. However, its performance starts to degrade on larger instances, evident from the largest case, where the gap spikes to around 3% and it hits the time limit. DRAGON performs favorably across all instances. While its gap is generally slightly higher than CP-SAT on small cases, it remains consistently below 0.5%, even on large-scale instances. In terms of trade-offs between performance and efficiency, DRAGON offers strong scalability and competitive quality, delivering high-quality solutions within reasonable reasoning time, making it highly effective for large-scale MKPs where traditional solvers may struggle.
4.5. Limitations and Future Directions
While DRAGON demonstrates strong potential in solving large-scale COPs, several limitations remain to be addressed. The quality of the reconstructed global solution is sensitive to both the decomposition strategy and the prompt formulation. Inappropriate segmentation or poorly structured prompts may lead to suboptimal or degenerated solutions, emphasizing the need for more robust and adaptive decomposition mechanisms. Moreover, the iterative nature of DRAGON introduces substantial computational overhead due to the large context required during each divide-and-conquer cycle, resulting in extended runtime. Although the current pipeline may not be efficient for small-scale problems, it proves particularly advantageous for super large-scale scenarios where traditional solvers and other learning-based methods struggle to deliver feasible or high-quality solutions across all testing tasks.
DRAGON also demonstrates that purely language-based agents can effectively address large-scale COPs, outperforming other prompt-based or code-generation-based approaches. Additionally, incorporating external optimization tools such as OR-Tools within the reconstruction stage enhances local refinement. However, the use of such tools often demands additional modeling expertise, especially for non-standard constrained sub-problem settings.
In the on-going work, we are integrating more external tools to automatically model customized constraints not only within the reconstruction agent but also into the decomposition process, with the goal of further reducing runtime and API costs while maintaining or improving solution quality. We also envision introducing a central coordination agent to better manage the interaction among decomposition and reconstruction agents, thereby improving the overall consistency and efficiency of the optimization process.
5. Conclusion
This paper presents DRAGON, a novel framework that leverages LLM agents to solve large-scale COPs via an iterative decomposition reconstruction process inspired by divide-and-conquer. Experiments across diverse COP domains (e.g., routing and packing) show that DRAGON consistently achieves strong performance, especially on extreme large instances where traditional solvers or basic prompt-based methods often fail due to scalability or timeouts. By decomposing complex problems into context manageable subproblems, DRAGON enables LLMs to reason effectively over intricate structures, identifying promising regions, and reconstructing feasible global solutions. This training-free, modular, and solver-agnostic approach supports flexible constraint handling and adapts easily to diverse COPs without specialized algorithm design. Although not optimized for speed, DRAGON avoids expensive code search and offers a strong balance between solution quality, adaptability, and cost, making it particularly well-suited for dynamic and large-scale settings. Future work will focus on enhancing decomposition strategies, improving prompt robustness, and integrating hybrid solvers to explore promising regions more efficiently.
Overall, DRAGON establishes a solid foundation toward general-purpose, interpretable, and scalable LLM-based optimization.
{acks}
This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RPGV-2025-017, AISG3-RP-2025-036-USNSF).
Appendix A Prompt Designs
A.1. General DRAGON Prompts
To implement DRAGON, we detail the design of the prompts used by the framework, which consists of two main agents: the Decomposer and the Reconstructor. These LLM prompts are designed to be general-purpose and adaptable across various combinatorial optimization problems (COPs). They contain placeholders that should be customized based on the specific COP being addressed.
A.1.1. Decomposer’s System Prompt
Decomposer System Prompt You are an expert in optimization with smart heuristics. The user will provide you an initial solution, whose data are formatted as {METADATA_FORMAT}. Help the user analysis the initial solution and point out which part is yet to be optimized using any creative heuristic methods you can. Ensure to output the answer in the specified required format. The problem inputs are: {INPUT_DATA} **Format**: Return only the node index, enclosed in <sub> and </sub>, separated by commas. For example: <sub>1,2,3</sub>
The placeholders above should be filled as follows:
  • METADATA_FORMAT: A list of key-value pairs describing the structure of the solution input.
  • INPUT_DATA: The actual raw input data to be passed to the LLM.
To align with our implementation, we recommend structuring the input data in a standardized JSON format, as illustrated in Table A1.
Table A1. Required keys in JSON format for different COPs. ✓indicates mandatory fields, while marks optional fields that are applicable in routing problems with additional constraints.
Key TSP CVRP BPP MKP
“name”
“type”
“num”
“depot”
“x”
“y”
“weights”
“values”
“capacity”
“demand”
“link”
Value types in the JSON schema are:
  • String (e.g., “name”, “type”),
  • Integer (e.g., “depot” for CVRP, “capacity” for BPP, “num” refer to node size for TSP/CVRP and item size for BPP/MKP),
  • List of integers (e.g., “x”, “y” for routing, “demand”, “capacity” for MKP).
For example of CVRP metadata:
Example JSON for CVRP
  "name": "CVRP-Example-001",
  "type": "cvrp",
  "num": 6,
  "depot": 0,
  "x": [50, 20, 40, 60, 80, 30],
  "y": [50, 70, 60, 40, 30, 90],
  "capacity": 100,
  "demand": [0, 10, 20, 30, 25, 15],
  "link": [
       [0, 5],
       [4, 2]
Decomposer’s User Prompt:
Decomposer User Prompt For the given optimization problem ({PROBLEM_TYPE}), my current solution is: {SOLUTION} These {SOLUTION_MARK} are not optimal. I want to improve them by remove some {ELEMENT} from the current solution, reconstruct them optimally. Please identify no more than {NUM} {ELEMENT} that could be improved.
Here’s how to define the placeholders:
  • PROBLEM_TYPE: Set as “tsp”, “cvrp”, “bpp”, or “mkp”.
  • SOLUTION: The current solution in a structured format (e.g., XML or JSON). For CVRP, the XML looks like:
Example XML for CVRP <sol> <route>0,2,3,0</route> <route>0,1,5,4,0</route> </sol>
Table A2. Mapping of user prompt placeholders for COPs.
Placeholder TSP CVRP BPP MKP
SOLUTION_MARK “route” “route” “pack” “pack”
ELEMENT “node” “node” “item” “item”
We set NUM = 20 in all experiments. We recommend not exceeding 50, as the identified elements will be passed to the reconstructor, which may face difficulty handling larger inputs.
A.1.2. Reconstructor’s System Prompt
Reconstructor System Prompt Act as an expert in combinatorial optimization. Your goal is find the best solution with given constraints, use any heuristic as you can. **Goal: Find the best solution, avoid the infeasible solutions.** **Requirements:** - {PROBLEM_SPECIFIC_REQUIREMENTS} - Check the feasibility of your solution to ensure the above conditions. OUTPUT FORMAT: After your solving, present your answer as: {SOLUTION_FORMAT}
Where the PROBLEM_SPECIFIC_REQUIREMENTS are:
  • TSP:
    • Must visit all the points exactly once, except the depot.
    • The start and end points must remain fixed.
    • The fixed path prevents any other visits in between; allow reversing of the fixed path where necessary.
  • CVRP:
    • Must visit all the customer nodes exactly once.
    • Each route must start and end at the depot.
    • Vehicle capacity must not be exceeded.
    • The fixed path prevents any other visits in between; allow reversing of the fixed path where necessary.
  • BPP:
    • Packed items’ total weight in each bin must be less than or equal to the bin’s capacity.
  • MKP:
    • Selected items’ total weight must be less than or equal to the knapsack’s capacity.
And the SOLUTION_FORMAT should follow:
  • TSP:
    <sol> <route>0,1,2,...,0</route> </sol>
  • CVRP:
    <sol> <route>0,1,2,...,0</route> <route>0,3,4,...,0</route> ... </sol>
  • BPP:
    <sol> <bin_0>0,1,2,...</bin_0> <bin_1>3,4,...</bin_1> ... </sol> Where bin_i is the i-th bin.
  • MKP:
    <sol> <knapsack_0>0,1,2,...</knapsack_0> <knapsack_1>3,4,...</knapsack_1> ... </sol> Where knapsack_i is the i-th knapsack.
Reconstructor’s User Prompt:
Reconstructor User Prompt Given a set of {ELEMENTS}, your task is to find the {OPTIMIZATION_OBJECT} while respecting the following constraints. {CONSTRAINTS} Current solution: {SOLUTION}
Where ELEMENT is set as:
  • “nodes” for TSP and CVRP
  • “items” for BPP and MKP
The OPTIMIZATION_OBJECT is set as:
  • TSP: “Find the shortest possible tour that visits all nodes exactly once and returns to the starting depot.”
  • CVRP: “Design a set of routes to deliver all customer demands using vehicles with limited capacity, minimizing the total distance while visiting each customer exactly once.”
  • BPP: “Given a set of items with their weights, your task is to find the best packing solution that minimizes the number of bins used while respecting each bin’s capacity.”
  • MKP: “Given a set of items with their values and weights, your task is to find the best packing solution that maximizes the total value while respecting each knapsack’s capacity.”
The CONSTRAINTS are:
  • Routing problem (TSP and CVRP):
    • “Fixed visiting path as (1,3), (5,9), …” (i.e., some segments in the tour must be preserved).
  • Packing problem (BPP and MKP):
    • Sum all items in the static segment $\mathbf{s}$ of each bin or knapsack, and treat them as a new bulky item, and set its weight as the sum of item weights. For MKP, further set the bulky item’s weight and value as the respective sums.
SOLUTION in positive experience storage should include the current solution in the same XML format as required by the SOLUTION_FORMAT in the system prompt. If you have collected infeasible solutions by previous rounds, you can continue to append your user prompt:
Reconstructor User Prompt (append revision) ... Analysis the following infeasible solution(s) and generate a new solution to meet the given constraints. {INFEASIBLE_SOL1} {INFEASIBLE_SOL2} {INFEASIBLE_SOL3} ...
Where the template for negative experience INFEASIBLE_SOLx can be:
<sol> ... </sol> <reason> Missing visit node(s): 3, 7 .... </reason>
Appendix B Detailed Experimental Results
B.1. TSPLIB Results
We then present full results on benchmarking instances from TSPLIB in Tables A3A7 add more commentary analysis
OPRO fails to return valid solutions for all groups, as some hard instances in groups exceed the runtime limit, resulting in an infi- nite gap inf. SGE consistently produces infeasible solutions (infe), likely because it lacks explicit feasibility checks, since its origi- nal design is for small problems (under 30 nodes). LMEA handles slightly larger instances but exhibits high gaps and fails beyond 2k-node due to prompt length limits and generation timeouts.
B.2. CVRPLIB Results
Full results of CVRPLIB instances are listed in Tables A8A9.
B.3. Ablation Studies
Ablation results analyzing decomposition and reconstruction strategies are provided in Table A10, and ablation results of LLM models are provided in Table A11.
B.4. Generalization Case Studies
BPP: Results are listed in Tables A12-A13. MKP: Performance are shown in Tables A14-A15.
Table A3. Result of DRAGON on the TSPLIB EUC_2D subset.
Size (k) Name Lower bound Objective value Gap (%) #API calls Input tokens (k) Output tokens (k) Running time (sec.)
<<0.1 berlin52 7542 8773 16.322 18 352.398 27.373 1153.372
eil51 426 440 3.286 25 355.136 33.297 1775.789
eil76 538 598 11.152 27 584.812 45.871 1678.698
pr76 108159 114421 5.790 16 293.463 20.841 929.784
rat99 1211 1402 15.772 18 257.759 26.649 1827.705
st70 675 757 12.148 24 409.572 35.584 1757.166
0.1-0.2 bier127 118282 135482 14.542 29 330.866 34.255 1957.364
ch130 6110 6503 6.432 32 533.508 60.685 2831.500
ch150 6528 7185 10.064 37 585.292 56.462 2719.040
d198 15780 17621 11.667 46 871.951 67.258 3499.006
eil101 629 694 10.334 40 705.357 70.520 2621.478
kroA100 21282 23160 8.824 40 690.147 59.315 2354.898
kroA150 26524 28278 6.613 36 502.292 47.662 1928.851
kroB100 22141 23417 5.763 40 507.247 64.044 3507.437
kroB150 26130 29267 12.005 9 223.708 17.918 613.975
kroC100 20749 22964 10.675 20 370.270 32.169 1566.927
kroD100 21294 22843 7.274 28 433.136 40.538 2048.876
kroE100 22068 23402 6.045 23 415.897 37.632 1696.224
lin105 14379 15979 11.127 17 198.105 22.608 1115.946
pr107 44303 44925 1.404 22 200.629 27.529 1556.138
pr124 59030 62151 5.287 22 386.629 35.707 1380.136
pr136 96772 101634 5.024 18 301.971 24.507 1008.675
pr144 58537 62220 6.292 18 347.802 25.506 932.202
pr152 73682 77233 4.819 13 202.484 14.085 745.549
rat195 $23$ 2588 11.408 26 438.158 47.778 2309.684
rd100 7910 8865 12.073 27 555.131 48.948 1984.560
u159 42080 47756 13.489 5 95.530 6.773 451.494
0.2-0.5 a280 2579 2901 12.485 19 358.893 27.183 988.452
d493 35002 39490 12.822 38 685.977 67.881 3444.888
fl417 11861 12776 7.714 22 350.584 28.860 1451.702
gil262 2378 2644 11.186 9 223.568 19.183 836.437
kroA200 29368 31920 8.690 31 603.447 45.735 2069.374
kroB200 29437 33157 12.637 40 803.297 74.733 3517.805
lin318 42029 46716 11.152 31 660.784 46.946 2264.698
pcb442 50778 58232 14.680 35 760.501 60.740 2542.813
pr226 80369 81657 1.603 19 336.121 23.377 807.854
pr264 49135 52991 7.848 29 574.093 49.003 1840.055
pr299 48191 53787 11.612 18 333.709 30.133 997.101
pr439 107217 121879 13.675 18 364.550 36.918 1166.819
rd400 15281 17215 12.656 13 261.578 20.461 1039.604
ts225 126643 144636 14.208 26 517.690 46.750 2419.392
tsp225 3919 4324 10.334 9 152.992 15.664 789.163
0.5-1 d657 48912 54577 11.582 26 618.878 54.094 1999.693
p654 34643 37476 8.178 19 518.997 32.475 1282.118
rat575 6773 7590 12.063 49 1085.435 82.549 3511.544
rat783 8806 9968 13.196 44 956.776 95.982 3505.717
u574 36905 40922 10.885 28 779.953 40.176 1516.220
u724 41910 46740 11.525 16 342.380 18.580 982.001
1-2 d1291 50801 60404 18.903 40 1640.458 61.223 2530.448
d1655 62128 73834 18.842 14 456.675 32.196 1164.084
fl1400 20127 21521 6.926 33 920.564 98.208 3246.920
fl1577 22249 25526 14.729 5 69.703 10.569 243.459
nrw1379 56638 63780 12.610 13 671.878 32.245 1241.157
pcb1173 56892 65849 15.744 9 462.186 18.142 625.697
pr1002 259045 293971 13.483 41 1609.575 62.281 2218.305
rl1304 252948 302389 19.546 33 895.836 69.866 2343.211
rl1323 270199 325813 20.583 22 846.612 42.888 1383.468
rl1889 316536 376474 18.936 29 867.053 98.208 2627.414
u1060 224094 248657 10.961 16 349.302 33.205 986.517
u1432 152970 172925 13.045 22 1033.374 39.479 1447.544
u1817 57201 68175 19.185 18 798.746 42.234 1986.882
vm1084 239297 267949 11.973 2 58.211 6.195 169.439
vm1748 336556 380867 13.166 43 1271.237 88.938 3346.815
2-5 d2103 80450 99775 24.021 23 888.182 74.305 2711.948
fl3795 28772 34268 19.102 26 1275.895 89.607 2589.477
fnl4461 152566 207168 35.789 18 1289.366 34.595 1760.015
pcb3038 137694 158664 15.229 18 714.400 67.205 2618.376
pr2392 378032 430293 13.824 38 1663.085 69.731 3113.234
u2152 64253 77075 19.955 33 1061.294 108.541 3208.794
u2319 234256 252285 7.696 21 632.640 78.352 2206.402
5-10 rl5915 655530 688752 5.068 14 1385.583 67.902 1733.749
rl5934 556045 679642 22.228 18 3465.698 149.070 2340.457
$\geq$10 brd14051 469445 530915 13.094 24 2502.501 37.871 1950.459
d15112 1573152 1780905 13.206 26 3877.022 49.884 2741.064
d18512 645488 733860 13.691 2 525.051 2.558 1012.969
rl11849 923368 1108759 20.078 12 1261.444 36.615 1398.704
usa13509 19982889 22827898 14.237 22 3515.226 117.270 2748.211
Table A4. Result of LMEA on the TSPLIB EUC_2D subset.
Size (k) Name Lower bound Objective value Gap (%) #API calls Input tokens (k) Output tokens (k) Running time (sec.)
<<0.1 berlin52 7542 22985 204.760 16 48.560 109.932 1633.713
eil51 426 1194 180.282 24 71.568 174.415 1953.738
eil76 538 1828 239.777 29 114.028 272.653 2818.638
pr76 108159 396874 266.936 37 150.923 346.013 3486.687
rat99 1211 6349 424.277 19 91.314 237.203 3511.408
st70 675 2710 301.481 33 122.232 289.233 3575.126
0.1-0.2 bier127 118282 522377 341.637 23 140.829 306.311 3585.385
ch130 6110 38924 537.054 22 131.648 226.152 3339.959
ch150 6528 48726 646.415 10 67.440 67.322 992.652
d198 15780 158386 903.714 18 160.992 232.834 2529.323
eil101 629 2813 347.218 16 78.112 139.750 1630.725
kroA100 21282 135707 537.661 26 129.116 272.577 3482.952
kroA150 26524 218343 723.190 6 41.580 80.523 1077.787
kroB100 22141 118419 434.840 35 174.020 412.730 3539.784
kroB150 26130 202969 676.766 24 166.296 200.905 3093.957
kroC100 20749 138488 567.444 16 79.440 174.127 2367.685
kroD100 21294 129635 508.787 17 84.303 154.815 2220.916
kroE100 22068 125387 468.185 35 174.020 413.782 3576.009
lin105 14379 102517 612.963 9 46.071 97.868 1092.099
pr107 44303 303855 585.856 22 117.128 285.457 3494.023
pr124 59030 526278 791.543 19 114.076 213.121 3426.667
pr136 96772 713073 636.859 14 90.776 163.846 1628.570
pr144 58537 118820 102.983 23 156.492 350.591 3294.159
pr152 73682 849178 1052.490 23 163.864 319.237 3442.378
rat195 $23$ 19502 739.518 21 177.534 294.969 3588.559
rd100 7910 44201 458.799 19 92.036 222.474 3590.330
u159 42080 347796 726.511 21 155.484 294.184 3547.076
0.2-0.5 a280 2579 30832 1095.502 19 221.996 172.781 2004.626
d493 35002 417749 1093.500 17 352.784 161.899 1966.156
fl417 11861 445256 3653.950 10 173.470 84.903 1325.056
gil262 2378 24519 931.077 8 89.032 79.375 1085.446
kroA200 29368 305520 940.316 16 142.192 168.055 2340.587
kroB200 29437 289238 882.566 26 231.062 297.838 3578.618
lin318 41345 544122 1194.635 8 108.504 89.533 837.863
pcb442 50778 723164 1324.168 8 147.848 56.720 606.033
pr226 80369 1442740 1695.145 21 212.100 198.219 1902.708
pr264 49135 999930 1935.067 14 162.673 109.308 1553.069
pr299 48191 677768 1306.420 23 299.092 272.404 3404.811
pr439 107217 1781733 1561.801 8 148.960 96.041 1393.436
rd400 15281 195457 1179.085 21 341.124 254.880 3520.224
ts225 126643 1339414 957.630 33 331.980 428.714 3322.474
tsp225 3919 30723 683.950 24 230.256 281.536 3326.143
0.5-1 d657 48912 816654 1569.639 8 218.376 78.842 810.460
p654 34643 1892175 5361.926 8 217.760 82.629 1254.455
rat575 6773 106143 1467.149 9 206.046 64.595 544.923
rat783 8806 167368 1800.613 11 338.778 101.534 1243.282
u574 36905 639997 1634.174 13 309.712 116.383 1127.479
u724 41910 814147 1842.608 13 387.374 125.961 1666.035
1-2 d1291 50801 1659285 3166.245 9 515.853 129.673 1757.716
d1655 62128 2123203 3317.466 10 782.590 106.926 976.751
fl1400 20127 1606464 7881.637 10 625.770 106.716 762.021
fl1577 22249 1310261 5789.078 8 584.680 125.307 1586.939
nrw1379 56638 1370252 2319.316 10 626.630 92.113 1529.937
pcb1173 56892 1362385 2294.686 8 402.384 114.889 1374.360
pr1002 259045 6148455 2273.508 12 494.088 140.352 1893.208
rl1304 252948 8924001 3427.998 12 700.392 147.164 1848.956
rl1323 270199 7697390 2748.786 8 474.632 96.158 864.331
rl1889 316536 14023995 4330.458 16 1466.704 214.204 2081.738
u1060 224094 6529359 2813.670 8 355.840 76.381 728.351
u1432 152970 3772731 2366.321 15 985.260 184.622 2175.513
u1817 57201 2040136 3466.609 10 870.070 133.765 1910.557
vm1084 239297 8130080 3297.485 13 594.997 168.779 1424.989
vm1748 336556 14404027 4179.831 17 1421.727 220.912 2143.461
2-5 d2103 80450 3136167 3798.281 8 828.416 95.846 974.011
fl3795 28772 inf inf - - - -
fnl4461 152566 inf inf - - - -
pcb3038 137694 inf inf - - - -
pr2392 378032 14868838 3833.222 9 1083.636 62.958 893.720
u2152 64253 2431659 3684.507 9 953.784 129.823 1924.544
u2319 234256 5867710 2404.828 9 1046.187 79.161 1211.242
5-10 rl5915 655530 inf inf - - - -
rl5934 556045 inf inf - - - -
$\geq$10 brd14051 469445 inf inf - - - -
d15112 1573152 inf inf - - - -
d18512 645488 inf inf - - - -
rl11849 923368 inf inf - - - -
usa13509 19982889 inf inf - - - -
Table A5. Result of OPRO on the TSPLIB EUC_2D subset.
Size (k) Name Lower bound Objective value Gap (%) #API calls Input tokens (k) Output tokens (k) Running time (sec.)
<<0.1 berlin52 7542 14774 95.899 9 12.420 2.428 61.642
eil51 426 730 71.369 12 17.760 3.098 111.630
eil76 538 1648 206.376 9 15.297 2.985 68.994
pr76 108159 277854 156.895 15 32.916 4.802 157.505
rat99 1211 3207 164.895 7 15.241 2.578 65.869
st70 675 inf inf - - - -
0.1-0.2 bier127 118282 inf inf - - - -
ch130 6110 inf inf - - - -
ch150 6528 inf inf - - - -
d198 15780 22512 42.664 9 41.007 4.905 65.386
eil101 629 inf inf - - - -
kroA100 21282 inf inf - - - -
kroA150 26524 inf inf - - - -
kroB100 22141 157184 609.926 12 27.366 4.655 75.213
kroB150 26130 273236 945.681 9 30.504 4.582 128.892
kroC100 20749 inf inf - - - -
kroD100 21294 127458 498.564 27 72.300 9.721 267.155
kroE100 22068 162325 635.569 9 21.474 3.301 96.389
lin105 14379 inf inf - - - -
pr107 44303 74143 67.354 9 22.896 3.671 73.170
pr124 59030 inf inf - - - -
pr136 96772 inf inf - - - -
pr144 58537 inf inf - - - -
pr152 73682 160979 118.479 9 31.986 4.099 119.534
rat195 $23$ inf inf - - - -
rd100 7910 inf inf - - - -
u159 42080 43375 3.079 9 33.348 4.138 77.461
0.2-0.5 a280 2579 2818 9.291 9 52.629 6.444 126.892
d493 35002 inf inf - - - -
fl417 11861 inf inf - - - -
gil262 2378 inf inf - - - -
kroA200 29368 328171 1017.444 7 29.741 4.716 80.113
kroB200 29437 327452 1012.384 9 40.176 5.487 164.678
lin318 41345 119866 185.200 7 48.956 4.542 83.120
pcb442 50778 542583 968.541 9 88.044 7.438 129.833
pr226 80369 110416 37.386 9 46.860 6.380 103.985
pr264 49135 77223 57.166 12 76.437 8.611 139.311
pr299 48191 83507 73.285 18 121.149 13.592 289.296
pr439 107217 inf inf - - - -
rd400 15281 215556 1310.617 9 74.589 6.980 173.999
ts225 126643 1484736 1072.380 15 81.471 9.050 145.865
tsp225 3919 inf inf - - - -
0.5-1 d657 48912 inf inf - - - -
p654 34643 inf inf - - - -
rat575 6773 inf inf - - - -
rat783 8806 inf inf - - - -
u574 36905 inf inf - - - -
u724 41910 inf inf - - - -
1-2 d1291 50801 inf inf - - - -
d1655 62128 inf inf - - - -
fl1400 20127 inf inf - - - -
fl1577 22249 inf inf - - - -
nrw1379 56638 inf inf - - - -
pcb1173 56892 inf inf - - - -
pr1002 259045 inf inf - - - -
rl1304 252948 inf inf - - - -
rl1323 270199 inf inf - - - -
rl1889 316536 inf inf - - - -
u1060 224094 inf inf - - - -
u1432 152970 inf inf - - - -
u1817 57201 inf inf - - - -
vm1084 239297 inf inf - - - -
vm1748 336556 inf inf - - - -
2-5 d2103 80450 inf inf - - - -
fl3795 28772 inf inf - - - -
fnl4461 152566 inf inf - - - -
pcb3038 137694 inf inf - - - -
pr2392 378032 inf inf - - - -
u2152 64253 inf inf - - - -
u2319 234256 inf inf - - - -
5-10 rl5915 655530 inf inf - - - -
rl5934 556045 inf inf - - - -
$\geq$10 brd14051 469445 inf inf - - - -
d15112 1573152 inf inf - - - -
d18512 645488 inf inf - - - -
rl11849 923368 inf inf - - - -
usa13509 19982889 inf inf - - - -
Table A6. Result of ReEvo(c) on the TSPLIB EUC_2D subset.
Size (k) Name Lower bound Objective value Gap (%) Inference time (sec.)
<<0.1 berlin52 7542 8608 14.140 0.034
eil51 426 453 6.470 0.133
eil76 538 579 7.620 0.190
pr76 108159 120153 11.090 0.214
rat99 1211 1361 12.410 0.199
st70 675 773 14.540 0.183
0.1-0.2 bier127 118282 131049 10.790 0.521
ch130 6110 6684 9.400 0.573
ch150 6528 7234 10.820 0.705
d198 15780 19247 21.980 1.614
eil101 629 690 9.840 0.305
kroA100 21282 22952 7.850 0.310
kroA150 26524 29605 11.620 0.812
kroB100 22141 24842 12.200 0.346
kroB150 26130 29532 13.020 0.812
kroC100 20749 24043 15.880 0.311
kroD100 21294 23910 12.290 0.341
kroE100 22068 23967 8.610 0.330
lin105 14379 15215 5.820 0.337
pr107 44303 47040 6.180 0.350
pr124 59030 68244 15.610 0.568
pr136 96772 108780 12.410 0.612
pr144 58537 65538 11.960 0.704
pr152 73682 83779 13.700 0.822
rat195 $23$ 2481 6.820 1.554
rd100 7910 8665 9.550 0.208
u159 42080 46487 10.470 0.946
0.2-0.5 a280 2579 3067 18.930 4.387
d493 35002 39701 13.430 22.934
fl417 11861 14132 19.150 14.108
gil262 2378 2685 12.940 3.699
kroA200 29368 32620 11.080 1.635
kroB200 29437 35024 18.980 1.687
lin318 41345 49017 16.630 6.110
pcb442 50778 57607 13.450 16.612
pr226 80369 94848 18.020 2.334
pr264 49135 57378 16.780 3.528
pr299 48191 58131 20.630 5.305
pr439 107217 127860 19.250 16.301
rd400 15281 17255 12.920 12.757
ts225 126643 134946 6.560 2.236
tsp225 3919 4350 11.020 2.319
0.5-1 d657 48912 56758 16.040 54.088
p654 34643 41121 18.700 53.232
rat575 6773 7638 12.780 36.041
rat783 8806 10018 13.770 89.825
u574 36905 44434 20.400 35.946
u724 41910 48979 16.870 72.058
1-2 d1291 50801 58731 15.610 402.530
d1655 62128 73008 17.510 850.869
fl1400 20127 24096 19.720 494.310
fl1577 22249 24944 12.110 734.272
nrw1379 56638 64780 14.380 497.545
pcb1173 56892 66840 17.490 308.277
pr1002 259045 310008 19.670 189.077
rl1304 252948 303725 20.070 421.773
rl1323 270199 317559 17.530 439.428
rl1889 316536 371919 17.500 1279.345
u1060 224094 266469 18.910 223.714
u1432 152970 172960 13.070 547.619
u1817 57201 66718 16.640 1120.653
vm1084 239297 280617 17.270 245.172
vm1748 336556 396194 17.720 1024.559
2-5 d2103 80450 86631 7.680 1763.146
fl3795 28772 inf inf 3600.101
fnl4461 152566 inf inf 3600.103
pcb3038 137694 inf inf 3600.029
pr2392 378032 459589 21.570 2617.235
u2152 64253 76497 19.060 1874.302
u2319 234256 248857 6.230 2318.946
5-10 rl5915 655530 inf inf 39401.084
rl5934 556045 inf inf 3600.103
$\geq$10 brd14051 469445 inf inf 3600.107
d15112 1573152 inf inf 3600.107
d18512 645488 inf inf 3600.109
rl11849 923368 inf inf 3600.096
usa13509 19982889 inf inf 3600.106
Table A7. Result of ReEvo(a) on the TSPLIB EUC_2D subset.
Size (k) Name Lower bound Objective value Gap (%) Inference time (sec.)
<<0.1 berlin52 7542 7549 0.100 1.873
eil51 426 444 4.410 1.199
eil76 538 561 4.310 1.276
pr76 108159 116960 8.140 3.068
rat99 1211 1287 6.350 3.670
st70 675 718 6.370 1.174
0.1-0.2 bier127 118282 125723 6.290 2.270
ch130 6110 6483 6.110 2.302
ch150 6528 $67$ 3.670 6.842
d198 15780 17404 10.290 3.839
eil101 629 669 6.380 1.743
kroA100 21282 22599 6.190 1.715
kroA150 26524 29047 9.510 2.709
kroB100 22141 23370 5.550 1.712
kroB150 26130 29227 11.860 2.752
kroC100 20749 21599 4.100 1.706
kroD100 21294 22831 7.220 1.703
kroE100 22068 23727 7.520 1.704
lin105 14379 14954 4.000 1.818
pr107 44303 47092 6.300 1.874
pr124 59030 61293 3.830 2.217
pr136 96772 108748 12.380 2.419
pr144 58537 60957 4.140 6.187
pr152 73682 76744 4.160 5.508
rat195 $23$ 2481 6.810 3.734
rd100 7910 8485 7.280 3.537
u159 42080 44710 6.250 5.823
0.2-0.5 a280 2579 2946 14.260 7.071
d493 35002 38921 11.200 12.795
fl417 11861 13568 14.400 9.877
gil262 2378 2618 10.100 10.935
kroA200 29368 32150 9.470 8.430
kroB200 29437 32924 11.850 8.610
lin318 41345 47067 11.990 17.050
lin318 42029 47067 11.990 17.050
pcb442 50778 59216 16.620 25.854
pr226 80369 89105 10.870 4.472
pr264 49135 54196 10.300 5.357
pr299 48191 53895 11.840 6.433
pr439 107217 119854 11.790 22.855
rd400 15281 17483 14.410 23.821
ts225 126643 132257 4.430 4.455
tsp225 3919 4306 9.880 9.928
0.5-1 d657 48912 56706 15.940 19.410
p654 34643 41239 19.040 40.870
rat575 6773 7789 15.000 15.848
rat783 8806 10293 16.890 25.583
u574 36905 43382 17.550 15.757
u724 41910 48290 15.220 22.880
1-2 d1291 50801 58520 15.200 70.189
d1655 62128 73286 17.960 161.547
fl1400 20127 24131 19.900 50.081
fl1577 22249 25637 15.230 134.016
nrw1379 56638 67191 18.630 52.296
pcb1173 56892 68553 20.500 91.328
pr1002 259045 303407 17.130 80.787
rl1304 252948 302084 19.430 94.308
rl1323 270199 316894 17.280 48.574
rl1889 316536 375685 18.690 86.597
u1060 224094 269036 20.060 41.572
u1432 152970 182777 19.490 107.057
u1817 57201 65498 14.510 199.884
vm1084 239297 283420 18.440 111.560
vm1748 336556 406986 20.930 171.830
2-5 d2103 80450 85620 6.430 103.428
fl3795 28772 34390 19.530 212.787
fnl4461 152566 221155 44.960 259.595
pcb3038 137694 166550 20.960 159.196
pr2392 378032 458366 21.250 247.613
u2152 64253 76332 18.800 100.445
u2319 234256 272864 16.480 103.202
5-10 rl5915 655530 686839 4.780 1019.255
rl5934 556045 674165 21.240 473.464
$\geq$10 brd14051 469445 577012 22.910 6294.430
d15112 1573152 1943279 23.530 7639.086
d18512 645488 794641 23.110 6152.544
rl11849 923368 1130865 22.470 2086.743
usa13509 19982889 25269411 26.460 3008.846
Table A8. Result of DRAGON on the CVRPLIB (X/XML) subset.
Size (k) Name Lower bound Objective value Gap (%) #API calls Input tokens (k) Output tokens (k) Running time (sec.)
0.1-0.2 X-n110-k13 14971 19298 28.903 16 235.218 27.996 974.263
X-n162-k11 14138 17280 22.224 15 330.999 31.953 792.177
0.2-0.5 X-n223-k34 40437 52877 30.764 13 249.044 23.135 710.500
X-n284-k15 20226 26916 33.076 31 754.018 64.900 2080.547
X-n420-k130 107798 133902 24.216 52 1238.116 110.004 3557.632
X-n459-k26 24139 31597 30.896 32 1088.167 80.217 2263.910
0.5-1 X-n513-k21 24201 31918 31.887 39 1445.657 84.518 2477.514
X-n670-k130 146332 211859 44.780 24 1181.644 75.815 2227.067
X-n701-k44 81923 94156 14.932 6 251.993 15.387 454.379
X-n783-k48 72386 94986 31.222 30 1117.956 88.571 2530.259
X-n801-k40 73311 81258 10.840 26 971.698 53.939 1820.531
1-2 X-n1001-k43 72355 85873 18.683 48 1985.763 125.876 3510.914
XML1000_1335_01 63968 73845 15.441 16 489.379 27.625 825.463
XML1500_2114_01 99781 112163 12.409 24 1855.369 101.138 2936.169
2-5 XML2500_2145_01 112858 154727 37.099 23 1557.874 76.582 1948.524
XML3500_3335_01 284329 300053 5.530 49 3397.328 161.908 3482.792
XML4500_3233_01 602565 625500 3.806 32 3373.244 150.978 3480.450
$\geq$5 XML5000_2224_01 315739 337855 7.005 28 5393.904 204.343 3549.789
XML5000_1321_01 1466910 1566502 6.789 21 3137.194 91.763 2735.975
XML5000_3135_01 396487 418520 5.557 26 2331.798 111.974 3028.920
Table A9. Result of ReEvo(a) on the CVRPLIB (X/XML) subset.
Size (k) Name Lower bound Objective value Gap (%) Inference time (sec.)
0.1-0.2 X-n110-k13 14971 17758.55 18.620 3.293
X-n162-k11 14138 17581.21 24.350 5.086
0.2-0.5 X-n223-k34 40437 47539.29 17.560 8.326
X-n284-k15 20226 25702.50 27.080 11.032
X-n420-k130 107798 122008.00 13.180 22.199
X-n459-k26 24139 32910.35 36.340 21.046
0.5-1 X-n513-k21 24201 34441.10 42.310 24.540
X-n670-k130 146332 179082.02 22.380 43.652
X-n701-k44 81923 99286.02 21.190 40.841
X-n783-k48 72386 101879.89 40.750 49.255
X-n801-k40 73311 87669.11 19.590 49.812
1-2 X-n1001-k43 72335 94073.34 30.020 72.013
XML1000_1335_01 63968 80300.27 25.530 69.735
XML1500_2114_01 99781 120350.96 20.620 101.063
2-5 XML2500_2145_01 112858 164532.37 45.790 189.593
XML3500_3335_01 284329 322280.75 13.350 289.669
XML4500_3233_01 602565 643929.79 6.860 418.924
$\geq$5 XML5000_2224_01 315739 359751.70 13.940 489.946
XML5000_1321_01 1466910 1570524.76 7.063 567.689
XML5000_3135_01 396487 446867.32 12.710 470.686
Table A10. Ablation for Decomposition and Reconstruction strategies
Decomposition Reconstruction Name Gap (%) Input tokens (k) Output tokens (k) Time (sec)
Random Heuristic eil51 5.634 - - 0.508
pr299 12.83 - - 1.043
u1060 12.166 - - 3.013
brd14051 13.621 - - 124.368
Average 10.938 - - 32.233
Solver eil51 5.634 - - 1.118
pr299 8.688 - - 601.031
u1060 12.374 - - 602.940
brd14051 13.386 - - 724.978
Average 10.021 - - 482.517
LLM eil51 7.981 66.143 10.258 101.244
pr299 13.851 93.981 14.448 821.860
u1060 11.173 89.304 11.820 656.243
brd14051 13.196 276.588 25.373 1520.241
Average 11.550 131.504 15.475 774.870
Heuristic Heuristic eil51 4.695 - - 0.504
pr299 15.688 - - 1.056
u1060 14.009 - - 3.195
brd14051 13.13 - - 157.806
Average 11.881 - - 40.640
Solver eil51 5.399 - - 600.516
pr299 10.207 - - 701.056
u1060 11.194 - - 603.169
brd14051 12.942 - - 753.731
Average 9.936 - - 664.618
LLM eil51 5.869 249.725 29.895 1653.207
pr299 10.747 105.327 16.817 833.485
u1060 13.170 75.322 15.279 1653.207
brd14051 13.374 104.041 19.172 954.097
Average 10.790 133.604 20.291 1273.499
LLM Heuristic eil51 5.634 196.966 9.220 407.764
pr299 11.436 318.755 11.862 458.776
u1060 10.961 420.500 17.615 462.020
brd14051 13.094 1389.088 3.652 1081.917
Average 10.281 581.327 10.587 602.619
Solver eil51 5.634 188.621 10.222 535.234
pr299 8.246 440.071 19.319 709.121
u1060 10.961 352.112 28.451 642.717
brd14051 13.079 10609.444 116.235 3260.043
Average 9.480 2897.562 43.557 1286.779
LLM eil51 3.286 355.136 33.297 1775.789
pr299 11.612 333.709 30.133 997.101
u1060 10.961 349.302 33.205 986.517
brd14051 13.094 2502.501 37.871 1950.459
Average 9.738 885.162 33.627 1427.466
Table A11. Impact of different LLM models
LLM model Name Gap (%) #API calls Input tokens (k) Output tokens (k) Running time (sec.)
gpt-4o berlin52 16.322 18 352.398 27.373 1153.372
kroB200 12.637 40 803.297 74.733 3517.805
d2103 24.021 23 888.182 74.305 2711.948
rl5934 22.228 18 3465.698 149.070 2340.457
rl11849 20.078 12 1261.444 36.615 1398.704
Average 19.057 22.20 1354.204 72.419 2224.457
deepseek-reasoner berlin52 13.710 4 2.226 48.926 2431.104
kroB200 14.190 4 5.031 45.960 2423.599
d2103 23.052 2 22.289 25.028 2183.935
rl5934 21.466 4 129.526 65.429 3036.950
rl11849 inf (Exceeds the model’s max input limit)
Average inf - - - -
o3 berlin52 3.248 20 10.425 182.739 3422.398
kroB200 10.147 12 15.368 152.977 3343.023
d2103 22.933 10 112.905 115.582 3220.190
rl5934 21.672 14 456.244 105.146 2220.981
rl11849 20.192 12 781.328 107.737 2150.865
Average 15.638 13.60 275.254 132.836 2871.491
gpt-4.1 berlin52 6.457 12 7.011 13.566 289.156
kroB200 8.900 13 18.041 17.039 284.020
d2103 23.693 12 135.993 18.674 421.396
rl5934 22.134 15 522.483 13.861 349.562
rl11849 19.552 12 781.658 15.665 433.936
Average 16.147 12.80 293.037 15.761 355.614
Table A12. Result of DRAGON for Bin Packing Problem.
Name Objective value L2 lower bound Running time (sec.) Gap (%) # API calls Input tokens (k) Output tokens (k)
Weibull 5k/test_0 2020 2012 512.607 0.398 19 238.147 28.797
Weibull 5k/test_1 1984 1983 527.750 0.050 15 195.512 22.628
Weibull 5k/test_2 1990 1978 122.899 0.607 4 40.045 7.187
Weibull 5k/test_3 1992 1986 513.842 0.302 18 240.698 32.893
Weibull 5k/test_4 1986 1980 762.268 0.303 17 235.260 29.430
Table A13. Result of ReEvo(a) for Bin Packing Problem.
Name Objective value L2 lower bound Running time (sec.) Gap (%)
Weibull 5k/test_0 2079 2012 57.687 3.330
Weibull 5k/test_1 2052 1983 55.274 3.480
Weibull 5k/test_2 2050 1990 55.965 3.640
Weibull 5k/test_3 2053 1992 57.682 3.374
Weibull 5k/test_4 2049 1986 56.779 3.485
Table A14. Result of DRAGON for MKP.
Name Objective value Upper bound Gap (%) Running time (sec.) # API calls Input tokens (k) Output tokens (k)
mkp_393_5_0 5427 5455 0.513 345.003 13 20.439 12.424
mkp_565_10_0 11415 11478 0.549 275.914 12 27.345 11.482
mkp_3594_20_0 37690 37706 0.042 159.304 12 143.961 6.545
mkp_3851_20_0 37018 37058 0.108 223.993 13 153.656 8.525
mkp_6870_50_0 83826 83880 0.064 213.504 12 275.633 12.221
mkp_6922_50_0 80305 80416 0.138 434.103 19 282.751 28.851
mkp_8864_50_0 85152 85266 0.134 320.866 16 350.271 17.373
mkp_10729_80_0 127696 127756 0.047 244.321 15 430.355 13.235
mkp_12397_50_0 105562 105592 0.028 123.530 7 240.881 5.239
mkp_30018_100_0 235436 235483 0.020 167.235 12 1152.107 3.960
Table A15. Result of OR-Tools (CP-SAT) for MKP.
Name Objective value Upper bound Gap (%) Running time (sec.)
mkp_393_5_0 5455 5455 0.000 2.52
mkp_565_10_0 11476 11478 0.017 3600.69
mkp_3594_20_0 37706 37706 0.000 80.11
mkp_3851_20_0 37058 37058 0.000 144.89
mkp_6870_50_0 83872 83880 0.010 3619.34
mkp_6922_50_0 80415 80416 0.001 3607.14
mkp_8864_50_0 85261 85266 0.006 3608.45
mkp_10729_80_0 127723 127756 0.026 3614.47
mkp_12397_50_0 105584 105592 0.008 3610.41
mkp_30018_100_0 227828 235483 3.251 3611.22

### v2

@@
-DRAGON: LLM-Driven Decomposition and Reconstruction Agents for Large-Scale Combinatorial Optimization
+arXiv:2601.06502v2[cs.AI] 02 Mar 2026
 Abstract.
@@
 We propose DRAGON, which stands for Decomposition and Reconstruction Agents Guided OptimizatioN, a novel framework that combines the strengths of metaheuristic design and LLM reasoning.
-Starting from an initial global solution, DRAGON autonomously identifies regions with high optimization potential and strategically decompose large-scale COPs into manageable subproblems.
+DRAGON autonomously identifies regions in initial solution with high optimization potential and strategically decompose large-scale COPs into manageable subproblems.
 Each subproblem is then reformulated as a concise, localized optimization task and solved through targeted LLM prompting guided by accumulated experiences.
-Finally, the locally optimized solutions are systematically reintegrated into the original global context to yield a significantly improved overall outcome.
-By continuously interacting with the optimization environment and leveraging an adaptive experience memory, the agents iteratively learn from feedback, effectively coupling symbolic reasoning with heuristic search.
-Empirical results show that, unlike existing LLM-based solvers limited to small-scale instances, DRAGON consistently produces feasible solutions on TSPLIB, CVRPLIB, and Weibull-5k bin packing benchmarks, and achieves near-optimal results (0.16% gap) on knapsack problems with over 3M variables.
+Finally, the locally optimized solutions are systematically reintegrated into the global context to yield a significantly improved outcome.
+By continuously interacting with the optimization environment and experience memory, the agents iteratively learn from feedback, effectively coupling symbolic reasoning with heuristic search.
+Empirical results show that, DRAGON consistently produces feasible solutions on TSPLIB, CVRPLIB, and Weibull-5k bin packing benchmarks, and achieves near-optimal results (0.16% gap) on knapsack problems with over 3M variables.
 This work shows the potential of feedback-driven language agents as a new paradigm for generalizable and interpretable large-scale optimization.
 Key words and phrases:
-Combinatorial Optimization, Large-scale, Metaheuristics
-Zhuoyi Lin is the corresponding author.
+Agentic AI; Combinatorial Optimization; Metaheuristics
+doi:
+SBAY6258
 1. Introduction
 Figure 1.
-Illustration of DRAGON on a TSP instance, where metadata represent the coordinates of locations to visit, serving as the main environment for the agents.
-The decomposer identifies a suboptimal active segment {4,6,7,8} (grayed) from a global solution.
+Illustration of DRAGON on a TSP instance.
+The decomposer identifies a suboptimal segment {4,6,7,8} (grayed) from a global solution.
 The segment is locally refined and reintegrated by the reconstructor to improve the global tour.
-Combinatorial optimization problems (COPs), such as the Traveling Salesman Problem (TSP), Vehicle Routing Problem (VRP), and Knapsack Problem (KP), are notoriously challenging due to their NP-hard nature Ausiello et al.
-(2012).
-These COPs are typically addressed using exact algorithms or heuristic methods, which often require substantial manual design and parameter tuning Rokbani et al.
-(2021); Cacchiani et al.
-(2022); Ma et al.
-(2023); Lin et al.
-(2024); Feng et al.
-(2025), thereby limiting their scalability and adaptability to varying problem instances and domains.
-Large Language Models (LLMs) have led to a series of breakthroughs for various challenging problems in natural language processing and artificial intelligence, including language understanding, text generation, code synthesis, and reasoning tasks (Brown et al., 2020; Chowdhery et al., 2023; Achiam et al., 2024).
-Recent evidence reveals that LLMs have emerged as powerful tools capable of addressing a variety of COPs, leveraging their reasoning capabilities for direct solution generation Yang et al.
-(2024); Wang et al.
-(2024) or heuristic design Ye et al.
-(2025); Romera-Paredes et al.
-(2024); Liu et al.
-(2024b); Zhu et al.
-(2025).
+Combinatorial optimization problems (COPs), such as the Traveling Salesman Problem (TSP), Vehicle Routing Problem (VRP), and Knapsack Problem (KP), are notoriously challenging due to the NP-hard nature ausiello2012complexity.
+COPs are typically addressed using exact algorithms or heuristic methods, which often require substantial manual design and parameter tuning rokbani2021bi; cacchianiKnapsackProblemsOverview2022; ma2023learning; lin2024cross; feng2025lifelong, thus limiting scalability and adaptability to varying problem instances and domains.
+Large Language Models (LLMs) have led to a series of breakthroughs for various challenging problems, including language understanding, text generation, code synthesis, and reasoning tasks (brown2020language; chowdhery2023palm; achiam2023gpt).
+Recent evidence reveals that LLMs have emerged as powerful tools capable of addressing a variety of COPs, leveraging their reasoning capabilities for direct solution generation yang2024large; wang2024can or heuristic design ye2024reevo; romera2024mathematical; liu2024evolution; zhu2025bridging.
 Their inherent strengths in abstraction, generalization, and semantic understanding allow them to generate promising solutions even without explicit algorithmic instructions.
-However, despite these advances, the current capabilities of LLMs for directly generating solutions remain largely confined to relatively small-scale instances, such as the TSP with fewer than 30 nodes Yang et al.
-(2024); Iklassov et al.
-(2024); Liu et al.
-(2024a).
-As problem size and complexity grow, LLM-based solutions often deteriorate due to inherent limitations, including restricted context length, reduced logical coherence, and difficulty in representing combinatorial structures Iklassov et al.
-(2024); Zhao et al.
-(2025).
-These limitations significantly impede the practical deployment of LLM-driven methods in realistic large-scale applications, such as logistics, transportation, and supply chain management, where problems commonly involve hundreds to thousands of nodes Bengio et al.
-(2021).
+However, the current capabilities of LLMs for directly generating solutions remain largely confined to relatively small-scale instances, such as TSP with fewer than 30 nodes yang2024large; iklassov2024self; liu2024large.
+As problem size and complexity increase, LLM-based solutions often deteriorate due to restricted context length, reduced logical coherence, and difficulty in representing combinatorial structures iklassov2024self; zhao2025can.
+These limitations significantly impede the practical deployment of LLM-driven methods in real large-scale applications, such as logistics, transportation, and supply chain management, where problems commonly involve hundreds to thousands of nodes bengio2021machine.
 In this study, we aim to answer two research questions: (1) Can an LLM agent, without solving the full COP, identify and isolate regions of the solution that are likely suboptimal or have improvement potential?
@@
 Prompt-based LLM methods often fail on large-scale instances due to the absence of domain knowledge and the context length limitation, while code-generation approaches require extensive task-specific training to develop robust heuristics, leading to substantial computational costs and poor cross-domain transferability.
-Meanwhile, classical metaheuristics such as divide-and-conquer and large neighborhood search (LNS) Pisinger and Ropke (2018) exhibit strong scalability in large-scale COPs by iteratively decomposing and refining solutions, though they rely heavily on handcrafted heuristics and expert knowledge.
+Meanwhile, classical metaheuristics such as divide-and-conquer and large neighborhood search (LNS) pisinger2018large exhibit strong scalability in large-scale COPs by iteratively decomposing and refining solutions, though they rely heavily on handcrafted heuristics and expert knowledge.
 To address the gap, we propose DRAGON, a framework that transforms complex large COP solving into decomposition and reconstruction tasks guided by LLM agents.
@@
 Among these efforts, a prominent direction involves using LLMs directly as solution generators, where carefully designed prompts and enriched input information guide the models to produce viable solutions.
-Early research by Wang et al.
-(2024) provided evidence that LLMs could solve certain graph-based COPs, although this capability was strongly tied to the quality of the provided prompts.
+Early research by wang2024can provided evidence that LLMs could solve certain graph-based COPs, although this capability was strongly tied to the quality of the provided prompts.
 Subsequent work has sought to refine this process by incorporating contextual information.
-Examples include integrating graph structural and topological data Wasserkrug et al.
-(2024), existing heuristic solutions Iklassov et al.
-(2024), explicit relationships between solutions and objectives Yang et al.
-(2024), and visual representations of problems or solutions Huang et al.
-(2024); Elhenawy et al.
-(2024a).
+Examples include integrating graph structural and topological data wasserkrug2024large, existing heuristic solutions iklassov2024self, explicit relationships between solutions and objectives yang2024large, and visual representations of problems or solutions huang2024multimodal; elhenawy2024eyeballing.
 Another research pathway involves leveraging LLMs as components within structured algorithms or frameworks.
-For instance, Liu et al.
-(2024a) demonstrated the utility of LLMs in evolutionary computation by prompting them to perform essential algorithmic operations such as selection, crossover, and mutation.
-Similarly, Elhenawy et al.
-(2024b) deployed multiple LLMs as collaborative agents, each responsible for distinct roles such as initialization, critique, and evaluation within a predefined optimization pipeline.
+For instance, liu2024large demonstrated the utility of LLMs in evolutionary computation by prompting them to perform essential algorithmic operations such as selection, crossover, and mutation.
+Similarly, elhenawy2024visual deployed multiple LLMs as collaborative agents, each responsible for distinct roles such as initialization, critique, and evaluation within a predefined optimization pipeline.
 Despite these promising advancements, current capabilities for directly generating solutions via LLMs remain largely restricted to relatively small-scale problems, typically limited to instances such as TSP with fewer than 30 nodes.
-To circumvent this limitation, Iklassov et al.
-(2024) introduced a decomposition approach wherein an LLM assesses problem complexity autonomously.
+To circumvent this limitation, iklassov2024self introduced a decomposition approach wherein an LLM assesses problem complexity autonomously.
 If deemed difficult, the model recursively subdivides the problem into simpler, manageable subproblems.
 Nonetheless, its scalability remains limited.
-Beyond direct solution generation, other studies have explored leveraging LLMs for heuristic design Romera-Paredes et al.
-(2024); Liu et al.
-(2024b); Ye et al.
-(2025) and mathematical modeling Xiao et al.
-(2023); AhmadiTeshnizi et al.
-(2024); Jiang et al.
-(2025b).
+Beyond direct solution generation, other studies have explored leveraging LLMs for heuristic design romera2024mathematical; liu2024evolution; ye2024reevo and mathematical modeling xiao2023chain; ahmaditeshnizioptimus; jiang2025droc.
 The former aims to use LLMs to discover heuristics expressed in code that have potential to outperform those crafted by human experts, while the latter focuses on translating natural language problem descriptions into mathematical formulations compatible with traditional OR solvers.
-More recently, research has investigated training large language models for end-to-end combinatorial optimization, rather than relying solely on prompting paradigms Jiang et al.
-(2025a, 2024).
-These directions fall outside the scope of this work, and we refer interested readers to Da Ros et al.
-(2025) for a comprehensive survey.
+More recently, research has investigated training large language models for end-to-end combinatorial optimization, rather than relying solely on prompting paradigms jianglarge; jiang2024bridge.
+These directions fall outside the scope of this work, and we refer interested readers to da2025large for a comprehensive survey.
 Large-Scale Combinatorial Optimization.
-Another research direction related to this work aiming to design metaheuristics for solving large-scale COPs, either with machine learning Li et al.
-(2021); Sun and Yang (2023) or through traditional methods Shaw (1998); Arnold et al.
-(2019).
+Another research direction related to this work aiming to design metaheuristics for solving large-scale COPs, either with machine learning li2021learning; sun2023difusco or through traditional methods shaw1998using; arnold2019efficiently.
 These approaches generally follow decomposition or divide-and-conquer principles.
@@
 These are compressed into a smaller subproblem; (c) A reconstruction step solves the reduced problem with additional constraints to yield a refined local solution.
-This is then reintegrated into the global solution.
-DRAGON is a two-stage framework that enables LLM agents to effectively solve large-scale COPs via state passing communication.
+DRAGON enables LLM agents to effectively solve large-scale COPs via state passing communication.
 As illustrated in Figure 2, DRAGON iteratively alternates between two core tasks: Decomposition and Reconstruction.
-Starting from an initial solution generated by any fast heuristic, DRAGON progressively refines it by identifying subregions with high improvement potential and solving localized subproblems under explicit constraint awareness.
-In each iteration, the current solution is decomposed into manageable subproblems, locally optimized through reconstruction agents, and then reassembled into a globally improved solution.
-This hierarchical process unifies local and global optimization, allowing DRAGON to scale from small instances to complex large-scale COPs.
-The example workflow illustrated in Figure 1 is summarized in Algorithm 1, where implicit communication between agents occurs via solution state passing, and the subsequent sections detail each component of the framework.
+Starting from fast heuristic initialization, DRAGON progressively refines it by identifying regions with high improvement potential and solving localized subproblems under explicit constraint awareness.
+In each iteration, the current solution is decomposed into manageable subproblems, locally optimized through reconstruction agents, and then reassembled into a globally improvement.
+This hierarchical process allows DRAGON to scale from small instances to complex large-scale COPs.
+Algorithm 1 summarized our workflow, where implicit communication between agents occurs via solution state passing.
 Algorithm 1 DRAGON for Large-Scale COPs
@@
 A central challenge in applying LLMs to large-scale COP decomposition lies in their limited capability to reason over complex global constraints.
-Although existing studies demonstrate the potential of LLMs in solving COPs Wilson et al.
-(2023); Yang et al.
-(2024), their capability to accurately detect and isolate sub-problems suitable for local refinement within large-scale COPs remains largely unexplored.
+Although existing studies demonstrate the potential of LLMs in solving COPs wilson2023; yang2024large, their capability to accurately detect and isolate sub-problems suitable for local refinement within large-scale COPs remains largely unexplored.
 To this end, we propose a decomposition agent ($\mathcal{D}$) which leverages LLMs to strategically partition COPs into smaller, manageable sub-problems for further refinement, rather than directly solving the entire COP.
@@
 It partitions the solution into active segments ($\mathbf{a}_{i}$), which have substantial improvement potential, and static segments ($\mathbf{s}_{i}$), which remain unchanged during the current iteration.
-Specifically, decomposition can be formulated as:
+Specifically, decomposition is formulated as:
 $(\mathbf{a}_{i},\mathbf{s}_{i})=\mathcal{D}(M,S_{i})$ (3)
@@
 The Compress step significantly reduces the size of the local data $|M_{i}^{\prime}|<<|M|$, making it manageable to downstream agents.
-However, it will introduce extra specific constraints, which are a set of compressed representations of static segment $\mathbf{s}_{i}$ (e.g., enforce edge preservation like $(9,5)\in C_{i}$).
+However, it will introduce extra specific constraints, which are a set of compressed representations of static segment $\mathbf{s}_{i}$.
 3.3. Reconstruction with Constraint Integration
-Recent studies Yang et al.
-(2024); Liu et al.
-(2024a); Iklassov et al.
-(2024) highlight the effectiveness of LLMs in solving small-scale COPs, typically with fewer than 30 decision variables.
+Recent studies yang2024large; liu2024large; iklassov2024self highlight the effectiveness of LLMs in solving small-scale COPs, typically with fewer than 30 decision variables.
 Building on this capability, our reconstruction agent ($\mathcal{R}$) is designed to solve compressed COP instances iteratively, derived from the decomposition stage.
@@
 To evaluate the effectiveness and generalizability of the proposed DRAGON framework, we conduct experiments on four representative COPs across routing, packing and assignment domains:
-  • Traveling Salesman Problem (TSP): A classical combinatorial routing problem that seeks the shortest possible tour visiting each city exactly once and returning to the starting point.
-  • Capacitated Vehicle Routing Problem (CVRP): Extension of TSP with multiple vehicles and demand constraints. The goal is to minimize total travel distance while satisfying delivery demands.
-  • Bin Packing Problem (BPP): Classical combinatorial optimization problem where items of various sizes must be packed into a minimal number of fixed-capacity bins without exceeding their limits.
-  • Multiple Knapsack Problem (MKP): A generalization of the 0–1 knapsack problem involving multiple independent knapsacks.
-Unlike the multi-dimensional knapsack problem Cacchiani et al.
-(2022), where each item is constrained by several resource limits, the MKP aims to maximize the total value by optimally assigning items to knapsacks without exceeding their individual capacities.
+$\bullet$ Traveling Salesman Problem (TSP): A classical combinatorial routing problem that seeks the shortest possible tour visiting each city exactly once and returning to the starting point.
+$\bullet$ Capacitated Vehicle Routing Problem (CVRP): Extension of TSP with multiple vehicles and demand constraints. The goal is to minimize total travel distance while satisfying delivery demands.
+$\bullet$ Bin Packing Problem (BPP): Classical combinatorial optimization problem where items of various sizes must be packed into a minimal number of fixed-capacity bins without exceeding their limits.
+$\bullet$ Multiple Knapsack Problem (MKP): A generalization of the 0–1 knapsack problem involving multiple independent knapsacks.
+Unlike the multi-dimensional knapsack problem cacchianiKnapsackProblemsOverview2022, where each item is constrained by several resource limits, the MKP aims to maximize the total value by optimally assigning items to knapsacks without exceeding their individual capacities.
 4.1.2. Datasets
-  • TSP We use 77 benchmark instances from TSPLIB Reinelt (1991) of the EUC_2D type, with sizes ranging from 50 to 20k nodes, covering a broad range of scales and spatial layouts.
-Oracle: optimal values or lower bounds reported in Reinelt (1991).
-  • CVRP We select subsets of 19 instances from CVRPLIB, including both X-type Uchoa et al.
-(2017) (up to 1000 nodes) and larger XML-type instances Queiroga et al.
-(2022) distribution, which was synthesized at large scale (up to 5k nodes) by Zhou et al.
-(2023).
-Oracle: optimal values or lower bounds reported in Uchoa et al.
-(2017); Zhou et al.
-(2023).
-  • BPP We adopt instances from FunSearch Romera-Paredes et al.
-(2024) and EoH Liu et al.
-(2024b) of all Weibull-5k, which is reflecting real-world scheduling and allocation scenarios.
-Oracle: L2 lower bounds determined by Romera-Paredes et al.
-(2024).
-  • MKP Following Google OR-Tools guidelines Perron and Furnon (2025), we generate 10 synthetic instances by with knapsack capacities uniformly sampled from $U(100,500)$, using 10 to 100 knapsacks.
+Detailed datasets for the aforementioned problems:
+$\bullet$ TSP We use 77 benchmark instances from TSPLIB reinelt1991tsplib of the EUC_2D type, with sizes ranging from 50 to 20k, covering a broad range of scales and spatial layouts.
+Oracle: optimal values or lower bounds reported in reinelt1991tsplib.
+$\bullet$ CVRP We select subsets of 19 instances from CVRPLIB, including both X-type UCHOA2017845 (up to 1k nodes) and larger XML-type instances queiroga2022 distribution, which was synthesized at large scale (up to 5k nodes) by zhou2023towards.
+Oracle: optimal values or lower bounds reported in UCHOA2017845; zhou2023towards.
+$\bullet$ BPP We adopt instances from FunSearch romera2024mathematical and EoH liu2024evolution of all Weibull-5k, which is reflecting real-world scheduling and allocation scenarios.
+Oracle: L2 lower bounds determined by romera2024mathematical.
+$\bullet$ MKP Following Google OR-Tools guidelines ortools, we generate 10 synthetic instances by with knapsack capacities uniformly sampled from $U(100,500)$, using 10 to 100 knapsacks.
 Item values and weights drawn uniformly from $U(1,100)$.
@@
 Prompting strategies for each stage of DRAGON are illustrated in Figure 2, with complete prompt templates provided in the Appendix (Prompt designs).
-We adopt simple global solution initialization algorithms to each problem: 1) Random Insertion Azar (1994) for TSP, 2) Greedy Nearest Neighbor heuristic Mohammed et al.
-(2017) for CVRP, and 3) First-Fit Decreasing Baker (1985) for both BPP and MKP.
-For comparative evaluation, we implement OR-Tools Perron and Furnon (2025) solvers as: CVRP is solved via the routing model Furnon and Perron (2025) with Guided_Local_Search [^1] https://developers.google.com/optimization/routing/routing_options as the metaheuristic, while BPP uses the CP-SAT Perron and Didier (2025).
+We adopt simple global solution initialization algorithms to each problem: 1) Random Insertion azar1994lower for TSP, 2) Greedy Nearest Neighbor heuristic mohammed2017solving for CVRP, and 3) First-Fit Decreasing baker1985new for both BPP and MKP.
+For comparative evaluation, we implement OR-Tools ortools solvers as: CVRP is solved via the routing model ortools_routing with Guided_Local_Search as the metaheuristic, while BPP uses the CP-SAT cpsatlp.
 All experiments run on an Ubuntu 20.04 server with an Intel Core i9-10900X (20 cores at 4.6 GHz) without GPU.
@@
 4.1.4. Feasibility check
-Feasibility checking for solutions obtained by any method is straightforward.
 We implement a checker based on the inherent constraints of each COP type.
@@
 DRAGON 25.56 29.35 26.73 15.51 15.48 6.45 20.15
-To evaluate performance on large-scale routing problems (TSP and CVRP), Table 1 presents the average optimality gap (%) across different instance size groups, comparing DRAGON with existing LLM-based solvers, including prompt-based methods (OPRO Yang et al.
-(2024), SGE Iklassov et al.
-(2024)) and code-generation approaches (LMEA Liu et al.
-(2024a), ReEvo Ye et al.
-(2025), with ReEvo(a) using Ant Colony Optimization (ACO) and ReEvo(c) employing a constructive heuristic).
+To evaluate performance on large-scale routing problems (TSP and CVRP), Table 1 presents the average optimality gap (%) across different instance size groups, comparing DRAGON with existing LLM-based solvers, including prompt-based methods (OPRO yang2024large, SGE iklassov2024self) and code-generation approaches (LMEA liu2024large, ReEvo ye2024reevo, with ReEvo(a) using Ant Colony Optimization (ACO) and ReEvo(c) employing a constructive heuristic).
 To ensure a fair comparison, all LLM-based methods use gpt-4o, with token usage and inference time recorded.
@@
 Figure 4. Average Optimality Gap (%) (log scale) across problem size groups for different methods on TSPLIB.
-As shown in Figure 4, results on TSPLIB reveal that SGE is excluded due to infeasible solutions, while OPRO is omitted as “inf” that no solution found due to its slow inference time across all groups.
+As shown in Figure 4, the results in TSPLIB reveal that SGE is excluded due to infeasibility, while OPRO is omitted as “inf” so that no solution was found due to its slow inference time in all groups.
 LMEA is able to solve up to medium-sized instances but exhibits relatively large optimality gaps.
 ReEvo(c) achieves strong results up to the medium-size group, whereas ReEvo(a) excels on smaller instances.
-Although DRAGON does not outperform across every group, it demonstrates clear superiority on larger-scale problems, benefiting from its task allocation mechanism, where decomposition and reconstruction collaborate effectively.
-Figure 5. Optimality Gap (%) for different methods on CVRPLIB.
+Although DRAGON is not the best on every group, it shows clear advantages on larger-scale problems, driven by its task‑allocation mechanism, where decomposition and reconstruction work effectively together.
+Figure 5. Optimality Gap (%) for methods on CVRPLIB.
 Figure 5 presents the methods capable of obtaining feasible solutions for CVRP instances.
 As CVRP is inherently more complex than TSP, DRAGON outperforms ReEvo(a) primarily on large-scale problems, where its decomposition–reconstruction mechanism demonstrates greater advantages.
-These findings highlight a major limitation of pure prompt-based methods: while capable of solving small to medium-scale COPs, their performance degrade significantly as problem size increases.
+These findings highlight a major limitation of pure prompt-based methods: their performance degrades significantly as problem size increases.
 This is mainly due to long context lengths and high decoding cost—making token-by-token generation is unstable when encounter large input loads.
@@
 Figure 7.
-Comparison of different LLM models by: (a) optimality gap, (b) number of API calls, (c) input tokens, (d) output tokens, (e) running time.
+Comparison of LLM models by: (a) optimality gap, (b) number of API calls, (c) input tokens, (d) output tokens, (e) running time.
 We evaluate general-purpose models (gpt-4o, gpt-4.1) and reasoning-oriented models (o3, r1).
-Gray bars indicate incomplete results due to token limit violations on the largest instance.
+Gray bars indicate incomplete results due to token limit violations.
 As shown in Figure 7, we evaluate the impact of different LLM backbones within the DRAGON framework, including general-purpose models (OpenAI gpt-4o, gpt-4.1) and reasoning models (OpenAI o3, DeepSeek r1).
@@
 For example, merging static segments ($\mathbf{s}$) into packed items requires less complex reasoning, allowing DRAGON to scale more effectively to larger instances.
-Results for EoH and EoH expert are reported directly from Liu et al.
-(2024b), where running time was not provided.
+Results for EoH and EoH expert are reported directly from liu2024evolution, where running time was not provided.
 While DRAGON achieves the lowest optimality gap, it is slower than code-generation based methods.
@@
 This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RPGV-2025-017, AISG3-RP-2025-036-USNSF).
+For Supplementary Materials: https://arxiv.org/abs/2601.06502
 Appendix A Prompt Designs
@@
 B.1. TSPLIB Results
-We then present full results on benchmarking instances from TSPLIB in Tables A3A7 add more commentary analysis
-OPRO fails to return valid solutions for all groups, as some hard instances in groups exceed the runtime limit, resulting in an infi- nite gap inf.
-SGE consistently produces infeasible solutions (infe), likely because it lacks explicit feasibility checks, since its origi- nal design is for small problems (under 30 nodes).
-LMEA handles slightly larger instances but exhibits high gaps and fails beyond 2k-node due to prompt length limits and generation timeouts.
+We then present full results on benchmarking instances from TSPLIB in Tables A3A7.
+OPRO fails to produce valid solutions across all groups, as several hard instances exceed the runtime limit, leading to infinite gaps (inf).
+SGE consistently returns infeasible solutions (infe), likely due to the absence of explicit feasibility checks, as it was originally designed for small-scale problems (fewer than 30 nodes).
+LMEA can handle moderately larger instances, but suffers from high optimality gaps and fails on instances beyond 2k nodes, primarily due to prompt length constraints and generation timeouts.
 B.2. CVRPLIB Results